OpenStudy (ammarah):

1a. Hyper V: (1,1) F(-5,1) Center: (2,1) 2b, Ellipse V: (1,2) (9,2) Covertices: (5,1)(5,3)

OpenStudy (ammarah):

@triciaal

OpenStudy (ammarah):

@reemii

OpenStudy (ammarah):

@satellite73

satellite73 (satellite73):

what?

satellite73 (satellite73):

general form for hyperbola is either \[\frac{x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] where the center is \((h,k)\) halway done already right?

OpenStudy (ammarah):

yeah

OpenStudy (ammarah):

(x-2)^2-(y-1)^2

OpenStudy (ammarah):

what are the denominators?

OpenStudy (ammarah):

the vertices squarede>

OpenStudy (ammarah):

@triciaal

OpenStudy (triciaal):

sorry was not here what is the question?

OpenStudy (triciaal):

did you substitute the values in the formula given by satellite73 to find the equation?

OpenStudy (ammarah):

Yeah

OpenStudy (ammarah):

I posted it above

OpenStudy (triciaal):

@ganeshie8

OpenStudy (ammarah):

I get no help seriously...